# FC–PINO: High-Precision Physics-Informed Neural Operators via Fourier Continuation

### v2

@@ Abstract
The physics-informed neural operator (PINO) is a machine learning paradigm that has demonstrated promising results for learning solutions to partial differential equations (PDEs). It leverages the Fourier Neural Operator (FNO) to learn solution operators in function spaces and leverages physics losses during training to penalize deviations from known physics laws. Spectral differentiation provides an efficient way to compute derivatives for the physics losses, but it inherently assumes periodicity. When applied to non-periodic functions, this assumption of periodicity can lead to significant errors, including Gibbs phenomena near domain boundaries which degrade the accuracy of both function representations and derivative computations, especially for higher order derivatives. To overcome this limitation, we introduce the FC–PINO (Fourier-Continuation-based Physics-Informed Neural Operator) architecture which extends the accuracy and efficiency of PINO and spectral differentiation to non-periodic and non-smooth PDEs. In FC–PINO, we propose integrating Fourier continuation into the PINO framework, and test two different continuation approaches: FC–Legendre and FC–Gram. By transforming non-periodic signals into periodic functions on extended domains in a well-conditioned manner, Fourier continuation enables fast and accurate derivative computations. This approach avoids the discretization sensitivity of finite differences and the memory overhead of automatic differentiation. We demonstrate that standard PINO fails (without padding) or struggles (even with padding) to solve non-periodic and non-smooth PDEs with high precision, across challenging benchmarks. In contrast, the proposed FC–PINO provides accurate, robust, and scalable solutions, substantially outperforming PINO alternatives, and demonstrating that Fourier continuation is critical for extending PINO to a wider range of PDE problems when high-precision solutions are needed.
@@ 1 Introduction
Partial differential equations (PDEs) are essential for modeling complex physical phenomena and dynamical systems in science and engineering. However, conventional numerical solvers become computationally intractable when applied to large-scale systems. To overcome these limitations, recent machine learning methods have been developed to learn solution operators of PDEs directly from data [1]. Among these, neural operators [2; 3; 4] have attracted considerable interest for their ability to process inputs defined on arbitrary grids and produce outputs at any spatial location. Moreover, they are supported by universal approximation guarantees for nonlinear operators [5].
@@
However, fully data-driven models can struggle in regimes with sparse or low-resolution data [6]. To mitigate this, prior knowledge of physics or conservation laws can be integrated into the learning process. Physics-Informed Neural Networks (PINNs) [7; 8; 9] are neural network representations of solutions to PDEs which can be learned by minimizing deviations away from physics laws. While PINNs are flexible and powerful, the associated optimization problem is often difficult, characterized by a highly nonconvex and non-smooth loss landscape [10; 11; 12; 13]. More recently, the Physics-Informed Neural Operator (PINO) [6] has been proposed and shown empirically to outperform PINNs on many tasks [14]. The original PINO relies on a Fourier Neural Operator (FNO) [2] which parametrizes the solution operator to a family of PDEs. Learning this operator in function space can be easier than the optimization problem arising when learning a single solution to a single PDE instance using PINNs. While the PINO framework was first proposed with a FNO as the backbone neural operator, it naturally extends to other neural operators such as graph neural operators [15].
@@
The accurate and efficient computation of derivatives is critical in PINO, as small numerical errors can be amplified through the physics losses and degrade solution quality. This is particularly critical in applications where high-precision solutions are required such as blowup modeling in fluid equations. Singularity formation, or blowups, for the Navier–Stokes equation poses one of the Clay Prize problems and is one of the most challenging problems in PDE analysis, with implications for applications in fluid modeling and beyond. While proving finite-time blowup for fluid equations under certain initial conditions remains a theoretical challenge, empirical approaches can play a valuable role in identifying potential blowup scenarios and approximate candidates for such events. If we can bound the error on these blowup candidates, then there is hope that (non)linear stability theory is sufficient to conclude that blowup does actually occur at a perturbation around the blowup candidate [16; 17; 18].
@@
One simple method of computing derivatives is finite differences, but it can require very fine resolutions to be accurate and can struggle with multi-scale or rapidly-varying dynamics. Automatic differentiation, widely used in PINNs [19], computes exact pointwise derivatives by applying the chain rule over a computational graph. It avoids the discretization error of finite differences but its memory cost can be prohibitive in large or deep models.
@@
On the other hand, spectral differentiation (also know as Fourier differentiation) offers a compelling alternative for smooth, periodic functions by computing derivatives in frequency space via simple multiplications of Fourier coefficients. It is accurate and has a cost of only $\mathcal{O}(N\log N)$ for $N$ grid points, with no increase in cost for higher derivatives. It is significantly more memory-efficient than automatic differentiation, especially for larger grids and larger models, as it avoids storing intermediate states. However, like other spectral methods, spectral differentiation implicitly assumes periodicity. When applied to non-periodic functions, it suffers from the Gibbs phenomenon which results in a loss of accuracy. While this loss of accuracy may be acceptable in certain contexts, it is a fundamental bottleneck when attempting to achieve high-precision solutions. This limitation motivates the development of Fourier-based techniques that remain highly accurate and robust for non-periodic inputs in physics-informed machine learning.
@@
One possibility is to use the Fourier series on a larger domain by padding or expanding the original domain. Such extension grants FNOs and PINOs the necessary expressivity to represent solutions for general non-periodic problems. However, a severe ill-conditioning issue can arise since the discrete Fourier transform is a Vandermonde matrix [20; 21; 22] if the extension is not done carefully. Zero-padding a signal on the sides introduces artificial discontinuities at the endpoints. When applying Fourier transforms to such signals, these discontinuities can result in Gibbs phenomena and spurious oscillations that do not reflect the true behavior of the underlying function (see Figure˜3 ). A common alternative to zero-padding is mirror padding, in which a signal is extended by reflecting it about one of its endpoints. This technique guarantees continuity of the signal at the mirror point, but introduces a discontinuity in the derivative unless the function is locally constant at that point. Consequently, the extended signal remains continuous but is not differentiable at the mirror point. Since the convergence rate of Fourier series depends on the regularity of the function being approximated, this lack of differentiability leads to very slow spectral convergence and persistent artifacts in the frequency domain representation.
@@
Ideally, to apply Fourier methods to problems with non-periodic solutions, one must convert the signal of interest to a periodic function in a well-conditioned manner. Fourier continuation (FC) methods [23; 24; 25] address this challenge by extending non-periodic signals to periodic ones over an enlarged domain through polynomial interpolation (see Figure˜1 for illustrations). These methods have been shown empirically to achieve superalgebraic convergence rates, making them particularly effective for spectral approximations of smooth, non-periodic data.
@@
Proposed Approach. We propose to use Fourier continuation (FC) as a mechanism that enables PINO to be applied to non-periodic functions with derivatives computed using spectral differentiation while retaining efficiency and high numerical precision. More precisely, we introduce the Fourier-Continuation-based Physics-Informed Neural Operator (FC–PINO), in which a FC layer is applied at the beginning of the FNO architecture. The FNO then operates on an extended domain, and its output is restricted immediately before the final projection operator. In FC–PINO, derivatives of the model output are computed using the chain rule by combining the analytic derivative of the projection operator with spectral differentiation on the extended domain of the pre-projection output. Because the FC layer enforces periodicity on the extended domain throughout the computation, spectral differentiation can be applied accurately and effectively before the restriction to the original domain. By construction, FC–PINO combines the universal operator approximation capabilities of neural operators with the convergence guarantees of FC, leading to errors in the solution derivatives that provably decay at an algebraic rate with increasing resolution. We also consider the use of two different FC methods as part of FC–PINO, the FC–Legendre and (Accelerated) FC–Gram [26] methods.
@@
Figure 1: Examples of periodic extensions obtained using the Fourier continuation methods. (top) Extensions of $f(x)=\sin(16x)-\cos(8x)$ from $[0,1]$ to [0.25,1.25][-0.25,1.25]. For display purposes only, a small constant was subtracted/added to the extended functions to prevent overlap. (bottom) Extensions of $f(x,y)=\sin(12x)-\cos(14y)+3xy$ from $0,1]^{2}$ to $0.25,1.25]^{2}$.
@@ 2.2.1 Physics-Informed Neural Networks (PINNs)
Physics-Informed Neural Networks (PINNs) [7; 8; 9] are neural networks designed to approximate solutions to PDEs. Their parameters are optimized by minimizing discrepancies not only from known solutions of the PDEs (when available) but also from physics constraints such as conservation laws and the governing PDEs. Over the past few years, a wide range of PINN variants have been developed and successfully applied to solve PDEs across various domains [27; 28; 29], including blowup problems [30; 31]. Notably, PINNs can be trained without relying on data, using only knowledge of the governing differential equations and of the associated boundary condition. Note that a PINN only learns the solution for a single PDE instance and does not generalize to new PDE instances without further optimization.
@@
In practice, the losses are evaluated on a discrete grid of points in the domain and summed up (possibly with quadrature weights) to approximate the loss integrals. The result is then used as the loss function for gradient-based optimization to train the model. The optimization could also be formulated via the variational form [32].
In these loss functions, $\alpha,\beta\in\mathbb{R}^{+}$ are constant hyperparameters used to balance the contributions of the different terms. Multi-objective optimization with physics losses can prove challenging with complex loss landscapes [33]. Gradients of the different losses can point in different directions with significant differences in magnitude, leading to unbalanced contributions and reductions of the different losses during training [34].
Manually tuning the loss coefficients can be computationally expensive, especially as the number of loss terms increases, but strategies have been developed to adaptively update these coefficients and mitigate this issue [35; 36; 37]. In this paper, we either conduct grid searches over the loss coefficients or employ ReLoBRaLo (Relative Loss Balancing with Random Lookback) [37] which uses the history of loss decay across random lookback intervals to achieve uniform progress across the different terms and faster convergence overall.
@@ 2.2.2 Physics-Informed Neural Operators (PINOs)
Physics-Informed Neural Operators (PINOs) [6] are closely related to PINNs, but rely on neural operators instead of neural networks to learn solution operators to entire families of PDEs simultaneously.
@@
 Neural networks have been successful in many machine learning applications but are fundamentally designed to map between finite-dimensional vectors, which can cause them to overfit to a specific discretization. In contrast, solving PDEs often involves working with functions and functional relationships defined over infinite-dimensional function spaces. Neural operators are a principled way to generalize neural networks to learn mappings between functions, and in particular, to learn solution operators of PDEs [3; 4]. They are applicable to functions given at any discretization and generate outputs that can be evaluated and show consistent accuracy across different resolutions. Neural operators also possess a universal approximation property for operators [38]. Neural operators have been successfully applied to a wide range of problems [39; 40; 41; 42; 43].
PINO relies on neural operators and typically on the Fourier Neural Operator (FNO) [2] which composes Fourier integral operator layers with pointwise nonlinear activation functions $\sigma$ to approximate non-linear operators as
@@
On uniform meshes, the Fourier transform $\mathcal{F}$ can be implemented using the fast Fourier transform. [44] maintain a comprehensive open-source library for learning neural operators in PyTorch, which serves as the foundation for our implementation. The FNO architecture is displayed in Figure 2.
Figure 2: The Fourier Neural Operator (FNO) architecture (extracted from [2]).
@@
 In the PINO framework [6], we can first learn a neural operator to approximate the groundtruth operator for a family of PDEs, and then fine-tune the model on any specific PDE instance. We refer to the operator learning phase as pretraining and the fine-tuning phase as instance fine-tuning or test-time optimization. In particular, with PINO, one can train from an unlimited amount of virtual PDE instances by drawing additional coefficient functions or initial conditions. Available data can also be used to supplement the training of the neural operator. A particularly helpful special case is to train a neural operator on low-resolution data instances with high-resolution PDE constraints. In this paper, we focus only on the more challenging setting where data is not available, although the entire exposition naturally applies to cases where data is available.
PINO has advantages over PINNs beyond the advantages of neural operators over neural networks, and has been shown to outperform PINNs on many PDE solving tasks [14]. Optimizing the coefficients in a basis function representation is easier than optimizing a single function as in PINNs. Also, learning these basis functions in the operator learning phase can make the later instance fine-tuning easier. In PINO, we do not need to propagate the information from the initial condition and boundary condition to the interior, it only requires fine-tuning the solution function parameterized by the solution operator.

### v3

@@ Abstract
The physics-informed neural operator (PINO) is a machine learning paradigm that has demonstrated promising results for learning solutions to partial differential equations (PDEs). It leverages the Fourier Neural Operator (FNO) to learn solution operators in function spaces and leverages physics losses during training to penalize deviations from known physics laws. Spectral differentiation provides an efficient way to compute derivatives for the physics losses, but it inherently assumes periodicity. When applied to non-periodic functions, this assumption of periodicity can lead to significant errors, including Gibbs phenomena near domain boundaries which degrade the accuracy of both function representations and derivative computations, especially for higher order derivatives. To overcome this limitation, we introduce the FC–PINO (Fourier‑Continuation-based Physics‑Informed Neural Operator) architecture which extends the accuracy and efficiency of PINO and spectral differentiation to non-periodic and non-smooth PDEs. In FC–PINO, we propose integrating Fourier continuation into the PINO framework, and test two different continuation approaches: FC–Legendre and FC–Gram. By transforming non-periodic signals into periodic functions on extended domains in a well-conditioned manner, Fourier continuation enables fast and accurate derivative computations. This approach avoids the discretization sensitivity of finite differences and the memory overhead of automatic differentiation. We demonstrate that standard PINO fails (without padding) or struggles (even with padding) to solve non-periodic and non-smooth PDEs with high precision, across challenging benchmarks. In contrast, the proposed FC–PINO provides accurate, robust, and scalable solutions, substantially outperforming PINO alternatives, and demonstrating that Fourier continuation is critical for extending PINO to a wider range of PDE problems when high-precision solutions are needed.
@@ 1 Introduction
Partial differential equations (PDEs) are essential for modeling complex physical phenomena and dynamical systems in science and engineering. However, conventional numerical solvers become computationally intractable when applied to large-scale systems. To overcome these limitations, recent machine learning methods have been developed to learn solution operators of PDEs directly from data (Kovachki et al., 2023). Among these, neural operators (Li et al., 2020; Azizzadenesheli et al., 2024; Berner et al., 2025) have attracted considerable interest for their ability to process inputs defined on arbitrary grids and produce outputs at any spatial location. Moreover, they are supported by universal approximation guarantees for nonlinear operators (Kovachki et al., 2021a).
@@
However, fully data-driven models can struggle in regimes with sparse or low-resolution data (Li et al., 2021). To mitigate this, prior knowledge of physics or conservation laws can be integrated into the learning process. Physics-Informed Neural Networks (PINNs) (Raissi et al., 2017a, b, 2019) are neural network representations of solutions to PDEs which can be learned by minimizing deviations away from physics laws. While PINNs are flexible and powerful, the associated optimization problem is often difficult, characterized by a highly nonconvex and non-smooth loss landscape (Wang et al., 2021b; Fuks and Tchelepi, 2020; Mattey and Ghosh, 2022; Krishnapriyan et al., 2021a). More recently, the Physics-Informed Neural Operator (PINO) (Li et al., 2021) has been proposed and shown empirically to outperform PINNs on many tasks (Rosofsky and Huerta, 2022). The original PINO relies on a Fourier Neural Operator (FNO) (Li et al., 2020; Duruisseaux et al., 2025) which parametrizes the solution operator to a family of PDEs. Learning this operator in function space can be easier than the optimization problem arising when learning a single solution to a single PDE instance using PINNs. While the PINO framework was first proposed with a FNO as the backbone neural operator, it naturally extends to other neural operators such as graph neural operators (Lin et al., 2025).
@@
The accurate and efficient computation of derivatives is critical in PINO, as small numerical errors can be amplified through the physics losses and degrade solution quality. This is particularly critical in applications where high-precision solutions are required such as blowup modeling in fluid equations. Singularity formation, or blowups, for the Navier–Stokes equation poses one of the Clay Prize problems and is one of the most challenging problems in PDE analysis, with implications for applications in fluid modeling and beyond. While proving finite-time blowup for fluid equations under certain initial conditions remains a theoretical challenge, empirical approaches can play a valuable role in identifying potential blowup scenarios and approximate candidates for such events. If we can bound the error on these blowup candidates, then there is hope that (non)linear stability theory is sufficient to conclude that blowup does actually occur at a perturbation around the blowup candidate (Chen et al., 2021; Chen and Hou, 2022; Wang et al., 2025a).
@@
One simple method of computing derivatives is finite differences, but it can require very fine resolutions to be accurate and can struggle with multi-scale or rapidly-varying dynamics. Automatic differentiation, widely used in PINNs (Baydin et al., 2017), computes exact pointwise derivatives by applying the chain rule over a computational graph. It avoids the discretization error of finite differences but its memory cost can be prohibitive in large or deep models.
@@
On the other hand, spectral differentiation (also know as Fourier differentiation) offers a compelling alternative for smooth, noise-free, periodic functions by computing derivatives in frequency space via simple multiplications of Fourier coefficients. It is accurate and has a cost of only $\mathcal{O}(N\log N)$ for $N$ grid points, with no increase in cost for higher derivatives. It is significantly more memory-efficient than automatic differentiation, especially for larger grids and larger models, as it avoids storing intermediate states.
However, like other spectral methods, spectral differentiation implicitly assumes periodicity. When applied to non-periodic functions, it suffers from the Gibbs phenomenon which results in a loss of accuracy. While this loss of accuracy may be acceptable in certain contexts, it is a fundamental bottleneck when attempting to achieve high-precision solutions. This limitation motivates the development of Fourier-based techniques that remain highly accurate and robust for non-periodic inputs in physics-informed machine learning.
@@
One possibility is to use the Fourier series on a larger domain by padding or expanding the original domain. Such extension grants FNOs and PINOs the necessary expressivity to represent solutions for general non-periodic problems. However, a severe ill-conditioning issue can arise since the discrete Fourier transform is a Vandermonde matrix (Higham, 2002; Press et al., 2007; Adcock et al., 2014) if the extension is not done carefully. Zero-padding a signal on the sides introduces artificial discontinuities at the endpoints. When applying Fourier transforms to such signals, these discontinuities can result in Gibbs phenomena and spurious oscillations that do not reflect the true behavior of the underlying function (see Figure˜3 ). A common alternative to zero-padding is mirror padding, in which a signal is extended by reflecting it about one of its endpoints. This technique guarantees continuity of the signal at the mirror point, but introduces a discontinuity in the derivative unless the function is locally constant at that point. Consequently, the extended signal remains continuous but is not differentiable at the mirror point. Since the convergence rate of Fourier series depends on the regularity of the function being approximated, this lack of differentiability leads to very slow spectral convergence and persistent artifacts in the frequency domain representation.
@@
Ideally, to apply Fourier methods to problems with non-periodic solutions, one must convert the signal of interest to a periodic function in a well-conditioned manner. Fourier continuation (FC) methods (Albin and Bruno, 2011; Fontana et al., 2020; Bruno and Paul, 2022) address this challenge by extending non-periodic signals to periodic ones over an enlarged domain through polynomial interpolation (see Figure˜1 for illustrations). These methods have been shown empirically to achieve superalgebraic convergence rates, making them particularly effective for spectral approximations of smooth, non-periodic data.
@@
Proposed Approach. We propose to use Fourier continuation (FC) as a mechanism that enables PINO to be applied to non-periodic functions with derivatives computed using spectral differentiation while retaining efficiency and high numerical precision. More precisely, we introduce the Fourier-Continuation-based Physics-Informed Neural Operator (FC–PINO), in which a FC layer is applied at the beginning of the FNO architecture. The FNO then operates on an extended domain, and its output is restricted immediately before the final projection operator. In FC–PINO, derivatives of the model output are computed using the chain rule by combining the analytic derivative of the projection operator with spectral differentiation on the extended domain of the pre-projection output. Because the FC layer enforces periodicity on the extended domain throughout the computation, spectral differentiation can be applied accurately and effectively before the restriction to the original domain. By construction, FC–PINO combines the universal operator approximation capabilities of neural operators with the convergence guarantees of FC, leading to errors in the solution derivatives that provably decay at an algebraic rate with increasing resolution. We also consider the use of two different FC methods as part of FC–PINO, the FC–Legendre and (Accelerated) FC–Gram (Amlani and Bruno, 2016) methods.
@@
Figure 1: Examples of periodic extensions obtained using the Fourier continuation methods. (top) Extensions of $f(x)=\sin(16x)-\cos(8x)$ from $[0,1]$ to [0.25,1.25][-0.25,1.25]. (bottom) Extensions of $f(x,y)=\sin(12x)-\cos(14y)+3xy$ from $0,1]^{2}$ to $0.25,1.25]^{2}$.
@@ 2.2.1 Physics-Informed Neural Networks (PINNs)
Physics-Informed Neural Networks (PINNs) (Raissi et al., 2017a, b, 2019) are neural networks designed to approximate solutions to PDEs. Their parameters are optimized by minimizing discrepancies not only from known solutions of the PDEs (when available) but also from physics constraints such as conservation laws and the governing PDEs. Over the past few years, a wide range of PINN variants have been developed and successfully applied to solve PDEs across various domains (Jagtap et al., 2020; Cai et al., 2022; Yu et al., 2022), including blowup problems (Wang et al., 2022, 2025b). Notably, PINNs can be trained without relying on data, using only knowledge of the governing differential equations and of the associated boundary condition. Note that a PINN only learns the solution for a single PDE instance and does not generalize to new PDE instances without further optimization.
@@
In practice, the losses are evaluated on a discrete grid of points in the domain and summed up (possibly with quadrature weights) to approximate the loss integrals. The result is then used as the loss function for gradient-based optimization to train the model. The optimization could also be formulated via the variational form (Weinan and Yu, 2018).
In these loss functions, $\alpha,\beta\in\mathbb{R}^{+}$ are constant hyperparameters used to balance the contributions of the different terms. Multi-objective optimization with physics losses can prove challenging with complex loss landscapes (Krishnapriyan et al., 2021b). Gradients of the different losses can point in different directions with significant differences in magnitude, leading to unbalanced contributions and reductions of the different losses during training (Wang et al., 2021a).
Manually tuning the loss coefficients can be computationally expensive, especially as the number of loss terms increases, but strategies have been developed to adaptively update these coefficients and mitigate this issue (Chen et al., 2018; Heydari et al., 2019; Bischof and Kraus, 2021). In this paper, we either conduct grid searches over the loss coefficients or employ ReLoBRaLo (Relative Loss Balancing with Random Lookback) (Bischof and Kraus, 2021) which uses the history of loss decay across random lookback intervals to achieve uniform progress across the different terms and faster convergence overall.
@@ 2.2.2 Physics-Informed Neural Operators (PINOs)
Physics-Informed Neural Operators (PINOs) (Li et al., 2021) are closely related to PINNs, but rely on neural operators instead of neural networks to learn solution operators to entire families of PDEs simultaneously.
@@
Neural networks have been successful in many machine learning applications but are fundamentally designed to map between finite-dimensional vectors, which can cause them to overfit to a specific discretization. In contrast, solving PDEs often involves working with functions and functional relationships defined over infinite-dimensional function spaces. Neural operators are a principled way to generalize neural networks to learn mappings between functions, and in particular, to learn solution operators of PDEs (Azizzadenesheli et al., 2024; Berner et al., 2025). They are applicable to functions given at any discretization and generate outputs that can be evaluated and show consistent accuracy across different resolutions. Neural operators also possess a universal approximation property for operators (Kovachki et al., 2021b). Neural operators have been successfully applied to a wide range of problems (Kurth et al., 2022; Gopakumar et al., 2023; Wen et al., 2023; Zhou et al., 2024; Ghafourpour et al., 2025).
PINO relies on neural operators and typically on the Fourier Neural Operator (FNO) (Li et al., 2020; Duruisseaux et al., 2025) which composes Fourier integral operator layers with pointwise nonlinear activation functions $\sigma$ to approximate non-linear operators as
@@
On uniform meshes, the Fourier transform $\mathcal{F}$ can be implemented using the fast Fourier transform. (Kossaifi et al., 2025) maintain a comprehensive open-source library for learning neural operators in PyTorch, which serves as the foundation for our implementation. The FNO architecture is displayed in Figure 2.
Figure 2: The Fourier Neural Operator (FNO) architecture (extracted from (Li et al., 2020)).
@@
In the PINO framework (Li et al., 2021), we can first learn a neural operator to approximate the groundtruth operator for a family of PDEs, and then fine-tune the model on any specific PDE instance. We refer to the operator learning phase as pretraining and the fine-tuning phase as instance fine-tuning or test-time optimization. In particular, with PINO, one can train from an unlimited amount of virtual PDE instances by drawing additional coefficient functions or initial conditions. Available data can also be used to supplement the training of the neural operator. A particularly helpful special case is to train a neural operator on low-resolution data instances with high-resolution PDE constraints. In this paper, we focus only on the more challenging setting where data is not available, although the entire exposition naturally applies to cases where data is available.
PINO has advantages over PINNs beyond those of neural operators over neural networks, and has been shown to outperform PINNs on many PDE solving tasks (Rosofsky and Huerta, 2022). Optimizing the coefficients in a basis function representation is easier than optimizing a single function as in PINNs. Also, learning these basis functions in the operator learning phase can make the later instance fine-tuning easier. In PINO, we do not need to propagate the information from the initial condition and boundary condition to the interior, it only requires fine-tuning the solution function parameterized by the solution operator.